Exercise 3. Suppose that |2 < 2. Prove that the series converges absolutely.
Answer #1

Let 2-2 , then to show 129 is convergent.

Note that ||+u - +ru = ud 8 ---------------------**

Now note that 1z|> 0 => 121 +n? > n => m2 - 2 + na

This gives us 5 + 5 11 + 2 | | 8 21

Since \{\sum \frac{8}{n^2}\} is convergent, by comparison test we get,  {Σεπ]}, thus \{\sum z_n\} converges absolutely.

Feel free to comment if you have any doubts. Cheers!